OpenStudy (anonymous):

whats x-2 over 5 equals 4x+2 over 25

OpenStudy (anonymous):

I think this is what you want \[\frac{x-2}{5} = \frac{4x+2}{25}\] ---------------------------------------------------------------------- If you want to solve for x, multiply by 25 on both sides to get \[\frac{25(x-2)}{5} = \frac{25(4x+2)}{25}\] Divide out common factors of the numerator and denominator on each side of the equation: \[5(x-2)=(4x+2)\] Multiply the left side of the equation out: \[5x-10 = 4x+2\] Subtract 4x from both sides: \[x-10 = 2\] Add 10 to both sides to get: \[x=12\]

OpenStudy (anonymous):

thank you so much(: , I've been struggling @uhhhhhh